# Acoustic transparency in dense granular suspensions

### v1

@@ Abstract
Surprisingly, no attempt has been made so far to interpret the acoustic response of compact sphere packings in terms of disorder correlations, even though such configurations were central to the original concept introduced by Torquato and collaborators [10]. Our interest in this question emerged unexpectedly during an experiment initially aimed at demonstrating that the diffusion coefficient of ultrasound in a dense granular suspension could be measured using a time-reversal experiment [11]. For this purpose, we employed a scattering sample composed of a random close-packed suspension of slightly polydisperse glass beads, with a diameter d=1.5± 0.1mmd=1.5\ $\pm\$0.1\ $\mathrm${mm}, and configured the setup such that the Time Reversal Mirror (TRM) was embedded deep within the scattering medium [see Fig. 1(a)]. In the first stage of the experiment, a short acoustic pulse was emitted by one element of a transducer array, and the resulting waveform was recorded by a receiver placed inside the suspension. As expected, the recorded signal featured an initial ballistic arrival followed by a long, multiply scattered coda [see Fig. 1(b)].
Unexpectedly, however, a closer examination revealed a striking feature: the frequency content of the ballistic pulse is significantly lower than that of the coda. This indicates that the medium exhibits strong dispersion (frequency-dependent attenuation caused by multiple scattering), resulting in a clear separation of the ballistic and scattered components not only in time, but also in frequency - a distinction clearly visible in the spectral domain [see Fig. 1(b)]. A comparison with the reference signal transmitted through water further highlights how propagation through the suspension enhances the low-frequency content of the ballistic wave relative to that of the incident pulse. A somewhat similar behavior was observed by Jia et al. in the late 1990s in dry granular media [12], though the underlying physics is fundamentally different: in dry packings, ultrasound propagates through the heterogeneous network of grain contacts. In a dense granular suspension, the acoustic wave travels primarily through the interstitial fluid (water), with weak coupling to the granular skeleton, as evidenced by the measured velocity of the ballistic wave, $c\sim 1700$ $\mathrm{m.s^{-1}}$, closely matching the speed of sound in water. The observation of this same spectral behavior in two physically distinct systems (a percolated solid skeleton in the first case, a pore fluid in the second) suggests that the effect may arise from structural disorder correlations inherent to the bead packing, which is the common feature shared by both media despite the different propagation mechanisms in each case.
@@
When the dilute medium is replaced by a random close packing of the same copper cylinders, a low-frequency ballistic pulse followed by a higher-frequency coda wave is observed, similar to the behavior seen in the 3D case [see left column of Fig. 3(b)]. As the sample thickness increases, high-frequency components tend to disappear, leaving only the low-frequency ballistic wave. As shown in the right column of Fig. 3(b), this progressive loss of high-frequencies results in a downward shift of the transmitted spectrum to about 0.3 MHz, bringing it closer to the lower edge of the bandwidth of the incident pulse. The remaining spectral components lie within a frequency range where scattering is significantly reduced, allowing the wave to propagate in an almost purely ballistic manner. This frequency range can therefore be referred to as a transparency window. This transparency window extends up to about 0.5 MHz, corresponding to $kd\sim 1.5$, where $k$ is the wavenumber in water. This lies well beyond the homogenization regime, characterized by $ks\ll 1$, where $s$ denotes the average spacing between cylinders, which is here is of the order of the cylinder diameter $d$; in that regime, transparency would be trivially expected. In other words, although the wave undergoes scattering, the resulting wave field fluctuations tend to cancel out due to disorder correlations. As in the 3D case, this suggests that disorder correlations must be taken into account to explain the observed behavior. The extinction length measured is 2.1 mm at 0.3 MHz , i.e., roughly three cylinder diameters, a surprisingly short distance for a frequency range described as a transparency window. However, this strong attenuation is likely due to absorption rather than scattering, and therefore does not contradict the notion of transparency in the sense of minimal wave scattering. We will return to this point later in the discussion.
@@
3D sphere packings - It has long been known that structural correlations can reduce (or enhance) scattering efficiency. This is, for example, the physical basis for the transparency of the cornea to visible light [14]. Such reductions in scattering efficiency can be quantified using an approach independently proposed by Saulnier [15] and Maret [16]. Given the form factor $F(q)$ of individual scatterers, which describes their response to an incident field, and the structure factor $S(q)$, which accounts for the spatial distribution of their positions, one can compute a corrected, or effective, scattering cross section, that incorporates interference effects between waves scattered by different scatterers :
@@
with $q$ the scattering wave number and $k_{r}$ the real part of the effective wave number [17].
Let us now evaluate the ability of this theoretical approach to account for the transparency window observed in the 3D glass bead packing. As a first step, we consider a suspension of glass beads (diameter $d=1.5\mathrm{mm}$) in water with a volume fraction $\Phi_{V}=40\%$. Although this specific case is not experimentally studied here, it is of particular interest because the structure factor can be derived analytically using the Percus - Yevick (PY) approximation [18]. The PY approximation applies to dense assemblies of monodisperse hard spheres, i.e., rigid particles with purely repulsive, short-range interactions. The resulting structure factor is shown in Fig. 4(a), either as a function of the scattering wavenumber (top panel), or as a function of both scattering angle and frequency (middle panel). Two main features emerge from the middle panel: first, a systematic reduction in scattering is observed in the vicinity of the forward direction (but not exactly along it). Second, a pronounced increase in scattering appears in the backward direction at frequencies around $qd=2\pi$, i.e., corresponding to a wavelength twice the glass bead diameter, reminiscent of the Bragg condition in ordered systems.
@@
Figure 4: Impact of structural correlations on ultrasound scattering in dense granular suspensions, for two packing fractions of glass beads. Top: structure factor $S(q)$ as a function of the reduced scattering wave number qdqd, where $q=\frac{4\pi}{\lambda}\sin(\theta/2)$, with $\lambda$ the wavelength in water, $\theta$ the scattering angle, and $d$ the bead diameter. The gray-shaded region highlights the low-$q$ range where $S(q)$ falls below unity (except at $q=0$ where $S(0)=1$ in any case). Middle: angular and frequency dependence of the structure factor. Bottom: Effective scattering cross section. Predictions based on the structure factor $S(q)$ (thick solid red line) and the GCPA model (thin solid blue line) show good agreement. In particular, both models predict a vanishing cross section within the frequency band marked by the gray-shaded area. For reference, the theoretical scattering cross section of a single glass bead [21] is shown as a dotted line.
@@
Figure 5: (a) Experimental setup used to measure acoustic transmission through a dense granular suspension of glass beads (diameter $d=1.5\mathrm{mm}$). The emitting transducer is excited with a half-period sine wave at its central frequency (3.5 MHz). The receiving transducer is mounted on a computer-controlled translation stage, allowing signal acquisition over source-receiver distances ranging from 5 mm to 50 mm in steps of 0.5 mm. (b) Top: transmitted signal amplitude as a function of frequency and propagation distance. Middle: spectra extracted from the top panel at two propagation distances and comparison with the spectrum in water (red solid line). Bottom: measured extinction length (crosses) compared with the GCPA prediction for the scattering mean free path (solid line).
The experimental setup is illustrated in Fig. 5(a). The sample was immersed in water between two identical ultrasonic transducers operating at a central frequency of 3.5 MHz. One transducer served as the emitter, driven by a half-period sine wave at its central frequency. The receiving transducer was mounted on a computer-controlled translation stage, allowing transmission measurements over a wide range of source-receiver distances (from 5 mm to 50 mm, with a step size of 0.5 mm). Fig. 5(b) (top and middle panels) shows how the transmitted signals evolve as a function of frequency and propagation distance. Remarkably, although the transducers are centered at a relatively high frequency, the combination of the broadband excitation and the very low attenuation within the transparency window allows the low-frequency components to dominate the transmitted signal. This makes it possible to clearly observe the transparency window predicted by the model, which extends up to $\sim$ 0.5 MHz (i.e., $kd\sim\pi$) in good agreement with the frequency range where the effective scattering cross section is predicted to vanish. In addition, a second transparency window is observed at higher frequencies, corresponding to the secondary dip in the predicted effective scattering cross section shown in the bottom panel of Fig. 4(b).
@@
In the transparency window, the frequency dependence of the corresponding absorption coefficient, defined as $\alpha_{\text{a}}=\frac{1}{2\ell_{\text{a}}}$, follows a power law: $\alpha_{a}=f^{\gamma}$, with $\gamma\approx 1.4$. This behavior departs significantly from standard theoretical models, which predict $\gamma=0.5$ for viscous dissipation localized in thin boundary layers between glass beads and the pore liquid (water) [23]. It also contrasts with earlier experimental studies, where exponents typically lie in the range $0.5<\gamma<1$ [24, 23].
Concluding remarks - Our main finding is the identification of an acoustic transparency window in glass bead packings immersed in water, allowing the propagation of a low-frequency ballistic wave excited by a broadband pulse emitted by a high-frequency transducer. We argued that this phenomenon can be interpreted as a consequence of spatial disorder correlations. To support this interpretation, we used an existing model that incorporates such correlations via the structure factor. We showed that its predictions agree well with those of the Generalized Coherent Potential Approximation (GCPA) model, which applies at high volume fractions, including the close-packing limit, but does not explicitly include disorder correlations. In the frequency range of the transparency window, unexplored in the seminal work by Page et al. [22], attenuation appears to be dominated by absorption rather than scattering so that the GCPA model, which does not incorporate absorption, fails to fit the experimental data. Moreover, the observed frequency dependence of absorption does not follow existing theoretical predictions, nor does it align with previous experimental results.
These observations highlight the need to revisit conventional models, which fail to account for absorption in dense granular suspensions. They also motivate further experimental investigation to better understand the mechanisms driving attenuation in such systems. To this end, we plan to expand our study by exploring a wider range of volume fractions, as well as varying bead materials and sizes.

### v2

# Acoustic transparency and absorption in dense granular suspensions

@@ Abstract
Surprisingly, no attempt has been made so far to interpret the acoustic response of compact sphere packings in terms of disorder correlations, even though such configurations were central to the original concept introduced by Torquato and collaborators [10]. Our interest in this question emerged unexpectedly during an experiment initially aimed at demonstrating that the diffusion coefficient of ultrasound in a dense granular suspension could be measured using a time-reversal experiment [11]. For this purpose, we employed a scattering sample composed of a random close-packed suspension of slightly polydisperse glass beads, with a diameter d=1.5± 0.1mmd=1.5\ $\pm\$0.1\ $\mathrm${mm}, and configured the setup such that the Time Reversal Mirror (TRM) was embedded deep within the scattering medium [see Fig. 1(a)]. In the first stage of the experiment, a short acoustic pulse was emitted by one element of a transducer array, and the resulting waveform was recorded by a receiver placed inside the suspension. As expected, the recorded signal featured an initial ballistic arrival followed by a long, multiply scattered coda [see the top panel of Fig. 1(b)].
Unexpectedly, however, a zoom-in on the first part of the signal revealed a striking feature: the central frequency of the ballistic pulse is significantly lower than that of the coda, clearly separating the ballistic and scattered components in both time and frequency. Comparison with the reference signal transmitted through water shows that propagation through the suspension enhances the low-frequency content of the ballistic wave relative to that of the incident pulse [see bottom panel of Fig. 1(b)]. We verified that the amplitude of this low-frequency ballistic wave scales linearly with the input pulse amplitude. In other words, the medium effectively acts as a linear low-pass filter. As discussed below, this effect arises from strong frequency-dependent scattering attenuation: the low-frequency components propagate essentially ballistically, whereas frequencies near the central frequency of the incident pulse undergo strong scattering.
A somewhat similar behavior was observed by Jia et al. in the late 1990s in dry granular media [12], though the underlying physics is fundamentally different: in dry granular packings, ultrasound propagates through the heterogeneous network of grain contacts. In a dense granular suspension, the acoustic wave travels primarily through the interstitial fluid (water), with weak coupling to the granular skeleton, as evidenced by the measured velocity of the ballistic wave, $c\sim 1700$ $\mathrm{m.s^{-1}}$, closely matching the speed of sound in water. It is consistent with the previous work by Brunet et al. [13]: adding a very small amount of liquid to a dry bead assembly leads to the formation of thin liquid films between the beads, resulting in such strong viscous dissipation that wave transmission through the solid skeleton (grain-contact network) is effectively suppressed.
The observation of this same spectral behavior in two physically distinct systems (a percolated solid skeleton in the first case, a pore fluid in the second) suggests that the effect may arise from structural disorder correlations inherent to the bead packing, which is the common feature shared by both media despite the different propagation mechanisms in each case.
@@
When the dilute medium is replaced by a random close packing of the same copper cylinders, a low-frequency ballistic pulse followed by a higher-frequency coda wave is observed, similar to the behavior seen in the 3D case [see left column of Fig. 3(b)]. As the sample thickness increases, high-frequency components tend to disappear, leaving only the low-frequency ballistic wave. As shown in the right column of Fig. 3(b), this progressive loss of high-frequencies results in a downward shift of the transmitted spectrum to about 0.3 MHz, bringing it closer to the lower edge of the bandwidth of the incident pulse. The remaining spectral components lie within a frequency range where scattering is significantly reduced, allowing the wave to propagate in an almost purely ballistic manner. This frequency range can therefore be referred to as a transparency window. This transparency window extends up to about 0.5 MHz, corresponding to $kd\sim 1.5$, where $k$ is the wavenumber in water. This lies well beyond the homogenization regime [15], characterized by $ks\ll 1$, where $s$ denotes the average spacing between cylinders, which is here is of the order of the cylinder diameter $d$; in that regime, transparency would be trivially expected. In other words, although the wave undergoes scattering, the resulting wave field fluctuations tend to cancel out due to disorder correlations. As in the 3D case, this suggests that disorder correlations must be taken into account to explain the observed behavior. The extinction length measured is 2.1 mm at 0.3 MHz , i.e., roughly three cylinder diameters, a surprisingly short distance for a frequency range described as a transparency window. However, this strong attenuation is likely due to absorption rather than scattering, and therefore does not contradict the notion of transparency in the sense of minimal wave scattering. We will return to this point later in the discussion.
@@
3D sphere packings - It has long been known that structural correlations can reduce (or enhance) scattering efficiency. This is, for example, the physical basis for the transparency of the cornea to visible light [16]. Such reductions in scattering efficiency can be quantified using an approach independently proposed by Saulnier [17] and Maret [18]. Given the form factor $F(q)$ of individual scatterers, which describes their response to an incident field as a function of the scattering wave number $q$, and the structure factor, defined as $S(q)=\frac{1}{N}\left|\sum_{j=1}^{N}e^{+i\textbf{q}\cdot\textbf{r}_{j}}\right|^{2}$, with $N$ the number of scatterers and $\textbf{r}_{i}$ their positions, one can compute a corrected, or effective, scattering cross section, that incorporates interference effects between waves scattered by different scatterers:
@@
with $k_{r}$ the real part of the effective wave number [19].
Let us now evaluate the ability of this theoretical approach to account for the transparency window observed in the 3D glass bead packing. As a first step, we consider a suspension of glass beads (diameter $d=1.5\mathrm{mm}$) in water with a volume fraction $\Phi_{V}=40\%$. Although this specific case is not experimentally studied here, it is of particular interest because the structure factor can be derived analytically using the Percus - Yevick (PY) approximation [20]. The PY approximation applies to dense assemblies of monodisperse hard spheres, i.e., rigid particles with purely repulsive, short-range interactions. The resulting structure factor is shown in Fig. 4(a), either as a function of the scattering wavenumber (top panel), or as a function of both scattering angle $\theta$ and frequency (middle panel) using $q=\frac{4\pi}{\lambda}\sin(\theta/2)$ with $\lambda$ the wavelength in water. Two main features emerge from the middle panel : first, the structure factor vanishes for all frequencies in the vicinity of the forward direction $\theta=0$ (but not exactly along it where it is equal to unity by definition), revealing a strong reduction in scattering thanks to Eq. 1. Second, in the backward direction $\theta=\pi$, the structure factor reaches a maximum at frequencies around $qd=2\pi$, i.e., corresponding to a wavelength twice the average distance (i.e., two bead diameters) between beads. This peak has an important contribution for the computation of the scattering cross-section, meaning that there is a strong backscattering. This is reminiscent of the Bragg condition in ordered system.
@@
Figure 4: Impact of structural correlations on ultrasound scattering in dense granular suspensions, for two packing fractions of glass beads. Top: structure factor $S(q)$, derived analytically using the Percus - Yevick approximation, as a function of the reduced scattering wave number $qd$, with $d$ the bead diameter. The gray-shaded region highlights the low-$q$ range where $S(q)$ falls below unity (except at $q=0$ where $S(0)=1$ in any case). Middle: structure factor plotted as a function of scattering angle and frequency, using $q=\frac{4\pi}{\lambda}\sin(\theta/2)$, where $\lambda$ is the wavelength in water and $\theta$ is the scattering angle. Bottom: effective scattering cross section. Predictions based on the structure factor $S(q)$ (thick solid red line) and the GCPA model (thin solid blue line) show good agreement. In particular, both models predict a vanishing cross section within the frequency band marked by the gray-shaded area. For reference, the theoretical scattering cross section of a single glass bead [23] is shown as a dotted line.
@@
Figure 5: (a) Experimental setup used to measure acoustic transmission through a dense granular suspension of glass beads (diameter $d=1.5\mathrm{mm}$). The emitting transducer is excited with a half-period sine wave at its central frequency (3.5 MHz). The receiving transducer is mounted on a computer-controlled translation stage, allowing signal acquisition over source-receiver distances ranging from 5 mm to 50 mm in steps of 0.5 mm. (b) Top: transmitted signal amplitude as a function of frequency and propagation distance. Middle: spectra extracted from the top panel at two propagation distances and comparison with the spectrum in water (red solid line). The three spectra are normalized to their respective maxima. Bottom: measured extinction length (crosses) compared with the GCPA prediction for the scattering mean free path (solid line). In the low frequency range, the corresponding extinction coefficient follows a power law (see inset).
The experimental setup is illustrated in Fig. 5(a). The sample was immersed in water between two identical ultrasonic transducers operating at a central frequency of 3.5 MHz. One transducer served as the emitter, driven by a half-period sine wave at its central frequency. The receiving transducer was mounted on a computer-controlled translation stage, allowing transmission measurements over a wide range of source-receiver distances (from 5 mm to 50 mm, with a step size of 0.5 mm). Fig. 5(b) (top and middle panels) shows how the transmitted signals evolve as a function of frequency and propagation distance. Remarkably, although the transducers are centered at a relatively high frequency, the combination of the broadband excitation and the very low attenuation within the transparency window allows the low-frequency components to dominate the transmitted signal. This makes it possible to clearly observe the transparency window predicted by the model, which extends up to $\sim$ 0.5 MHz (i.e., $kd\sim\pi$) in good agreement with the frequency range where the effective scattering cross section is predicted to vanish. Two other frequency ranges with reduced scattering are observed around 1.4 MHz and 3 MHz, corresponding to local maxima in the extinction length, with the first one consistent with the second dip observed at $d\sim 1.4\lambda$ in the scattering cross section shown in the bottom panel of Fig. 4(b).
@@
In the transparency window, the frequency dependence of the corresponding absorption coefficient, defined as $\alpha_{\text{a}}=\frac{1}{2\ell_{\text{a}}}$, follows a power law: $\alpha_{a}\sim f^{\gamma}$, with $\gamma\approx 1.4$ [see the inset in the bottom panel of Fig. 5(b)]. This behavior departs significantly from standard theoretical models, which predict $\gamma=0.5$ for viscous dissipation localized in thin boundary layers between glass beads and the pore liquid (water). For example, the model, as described in [25], predicts that the absorption coefficient is given by $\alpha_{a}=\Phi_{V}/(1-\Phi_{V})d^{-1}c^{-1}\sqrt{4\pi\mu f/\rho}$, with $\mu$ and $\rho$ the viscosity and mass density of water, respectively. The absorption length estimated from this model gives $\ell_{\text{a}}=400$ mm at 0.3 MHz corresponding to $260$ bead diameters. This value is more than one order of magnitude larger than the measured value $\ell_{\text{ext}}\sim\ell_{\text{a}}\approx 20~\text{mm}$. The measured values also contrast with earlier experimental studies, where exponents typically lie in the range $0.5<\gamma<1$ [26, 25].
Concluding remarks - Our main finding is the identification of an acoustic transparency window in dense granular suspensions, allowing the propagation of a low-frequency ballistic wave excited by a broadband pulse emitted from a high-frequency transducer. Although such experimental packings of slightly polydisperse glass beads immersed in water are unlikely to satisfy the conditions required for a hyperuniform medium, they clearly exhibit at least short-range correlations. We argued that the observed transparency window can be interpreted as a consequence of these correlations. To support this interpretation, we used an existing model that incorporates such correlations via the structure factor. We showed that its predictions agree well with those of the Generalized Coherent Potential Approximation (GCPA) model, which applies at high volume fractions, including the close-packing limit, but does not explicitly include disorder correlations. In the frequency range of the transparency window, unexplored in the seminal work by Page et al. [24], attenuation appears to be dominated by absorption rather than scattering so that the GCPA model, which does not incorporate absorption, fails to fit the experimental data. Moreover, the observed frequency dependence of absorption does not follow existing theoretical predictions, nor does it align with previous experimental results. These results motivate further experimental investigation to better understand the mechanisms driving attenuation in such systems. To this end, we plan to expand our study by exploring a wider range of volume fractions, as well as varying bead materials and sizes.